OpenStudy (anonymous):

What value corresponds to the 60th percentile in the following data set? {12, 28, 35, 42, 47, 49, 50}

OpenStudy (inkyvoyd):

There's 7 numbers in here, so you are looking for the 0.6*7 (round up) number in the series.

OpenStudy (amistre64):

just like a median divides up the set in half; "*tiles" divide the set up into constituent parts as well

OpenStudy (anonymous):

so its 49

OpenStudy (inkyvoyd):

47, actually. first you multiple the percentile by the total number of people, then if it is a whole number, take the average between that number in the series, and the next. If it is not a whole number, simply round up, and find that number in the series and the answer is there. In this case, 0.6*7=4.2 Thus we round up, because it is not a whole number. We get 5, then we find the fifth number, with the num bers placed in order and from left to right. If you are having trouble, try googling "wikipedia percentile"

OpenStudy (amistre64):

7/100 = 1% 6(.07 = 1%) 4.2 = 6% ; rnd up = 5th position

OpenStudy (anonymous):

percentile is p60=30.5+[(60%7-28)/(75))(8)=33.038